OpenStudy (anonymous):

Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. (Hint: A chord divides a circle into two segments. In problem 1, you found the area of the smaller segment.)

OpenStudy (anonymous):

find the smallest and substract it from phi*8^2

OpenStudy (anonymous):

the smallest one is 5.07 but idk how to do it

OpenStudy (anonymous):

tha way i told ya earlier...find the smallest ,subtract it and get the largest

OpenStudy (anonymous):

subratct it from what though?

OpenStudy (anonymous):

from phi*8^2

OpenStudy (anonymous):

my answer is suppost to be like this \[A = \left[\begin{matrix}? & ? \\ ? & ?\end{matrix}\right] \pi + __\sqrt{_} ) inches 2\]

OpenStudy (anonymous):